OpenStudy (anonymous):

derivative find dy/dx y=(x^3-2x^2)(3x+4)(x^4-5x^3+9)

OpenStudy (anonymous):

y = (3x^4 - 6x^3 + 4x^3 - 8x^2)(x^4 - 5x^3 + 9) y = (3x^4 - 2x^3 - 8x^2)(x^4 - 5x^3 + 9) y = (3x^8 - 2x^7 - 8x^6 - 15x^7 + 10x^6 + 40x^5 + 27x^4 - 18x^3 - 72x^2) y = 3x^8 - 17x^7 + 2x^6 + 40x^5 + 27x^4 - 18x^3 -72x^2 dy/dx = 24x^7 - 119x^6 + 12x^5 + 200x^4 + 108x^3 - 54x^2 - 144x